Jump to content

Problem with one channel


Guest guest
 Share

Recommended Posts

Hi everyone,

I seem to be having a problem with LOR. This problem occurs while a show is playing and when I use the hardware utility. One channel does not turn completely off. It looks like it is stuck at 25% on when it should be 0%. When I send a command either manually via hardware util or automatically through a show it comes on fine it just doesn't turn off all the way. Something else I've noticed is whenever I do a shimmer or twinkle effect on the channel whenever the effect is over a pulsing effect lingers for 10 or so seconds and then dies out. The problem first started when I plugged the box in. Even without being hooked up to a computer this one channel immediately started pulsing. Eventually it went out on it's own and worked fine for about a day and now it is acting up again.

I have a video of this behavior if it would help the LOR people diagnose.

Link to comment
Share on other sites

Sounds like it is a faulty triac or opto isolator... Send an email to support@lightorama.com and tell mention a stuck channel. Tell them the type of controller, shipping information for a replacement and if you have it, the orginal order number (last 4 digits is enough).

Link to comment
Share on other sites

Guest
This topic is now closed to further replies.
 Share

×
×
  • Create New...